# Embedding Self-Correction as an Inherent Ability in Large Language Models for Enhanced Mathematical Reasoning

### v1

@@ 3.2.2 CoSC Self Enhancement
Dense solution sampling. For mathematical questions, there are usually multiple solutions, but use GPT-4 for annotating them is expensive. Therefore, we use the seed model after CoSC foundational learning to resample the questions in the datasets multiple times further improve the model generalization ability in answering questions. Specifically, we use the CodeLLaMA-34B model which has the best performance after CoSC foundational learning to perform dense nucleus-sampling on 16k training data questions. Each question in the MATH and GSM8K training datasets is sampled 64 times. We filter out the correct answers based on whether they match the ground-truth.
Dense question sampling. Similarly, there are multiple ways to ask a mathematical question. To improve the model generalization ability of question understanding, we use the data from MetaMath (Yu et al., 2023) to generalize the questions. Each question in MetaMATH is rewrited by simply rephrasing it, self-verifying the conditions in the question, adding if-then questions to reverse the question conditions, etc. Specifically, we use the CodeLLaMA-34B model with CoSC foundational learning to perform nucleus-sampling on the questions, sampling each question once, and filtering out data points based on whether they match the ground-truth.
In the end, we obtain a total of 339k data points, including 37k seeding data generated from GPT-4 and 302k generated from the CodeLLaMA-34B model with CoSC foundational learning. Then, we adopt them to train models from scratch to obtain the final our CoSC model.
3.3 Implementation Details
By using 339k data points, including 37k seeding data and 302k self-generated data, we fine-tune the base models of LLaMA-2 (Touvron et al., 2023) and CodeLLaMA (Roziere et al., 2023) to obtain our CoSC and CoSC-Code, respectively. They are with different parameter sizes, such as 7B, 13B, and 34B. All models use full-scale fine-tuning. We use the AdamW optimizer with a learning rate of 2e-5 for all models, with a batch size set to 128, training for 1 epoch. To enable training, we use DeepSpeed ZeRO stage 3 (Rajbhandari et al., 2021) and Flash-Attention 2 (Dao, 2023) to optimize the model’s memory usage. During inference, we set a maximum of 3 calls to the Python interpreter and a maximum token length of 2048. Our experiments train the models in 7B size with 8 NVIDIA A800 80GB GPUs and train the models in 13B and 34B with 16 NVIDIA A800 80GB GPUs.
4 Experiments
4.1 Evaluation Setup
Datasets. We evaluated models on the most widely used mathematical problem tasks, MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021). The MATH dataset encompasses a total of 12,500 problems, partitioned into 7,500 for training and 5,000 for testing. The GSM8K dataset contains 8,500 problems, with 7,500 for training and 1,000 for testing. These datasets collectively encompass a broad range of mathematical questions, from basic arithmetic to competition level. More details of datasets are provided in Appendix C.
Table 2: Accuracy results (%) on MATH and GSM8K datasets. Vanilla models are tested with CoT. ZS indicates the zero-shot inference without demonstrations. PAL refers to the Program-Aided Language model prompting (Gao et al., 2023). The best results in each section are in bold and the second-best results are underlined. Our CoSC can outperform open-sourced LLMs, proprietary LLMs and some of multi-modal LLMs, on the more challenging MATH dataset.
@@ 4.1 Evaluation Setup
GPT-4o - - 76.6 - -
GPT-4V - - 52.9 92.0 72.5
GPT-4 (PAL) - - 51.8 94.2 73.0
GPT-4 - - 42.5 92.0 67.3
ChatGPT (PAL) - - 38.7 78.6 58.7
ChatGPT - - 35.5 80.8 58.2
Claude-3.5 Sonnet - - 71.1 96.4 83.8
Claude-3 Opus - - 60.1 95.0 77.6
Claude-2 - - 32.5 85.2 58.9
Gemini-1.5 Pro - - 67.7 90.8 79.3
Gemini-1.5 Flash - - 54.9 86.2 70.6
Gemini-1.0 Ultra - - 53.2 88.9 71.1
Gemini-1.0 Pro - - 32.6 77.9 55.3
PaLM-2 - 540B 34.3 80.7 57.5
@@
CodeLLaMA (PAL) CodeLLaMA 7B 16.6 34.0 25.3
Platypus-2 (Lee et al., 2023) LLaMA-2 7B 5.4 14.4 9.9
@@
CoSC-Code (Ours) CodeLLaMA 7B 47.6 74.7 61.2 (+2.6)
LLaMA-2 LLaMA-2 13B 6.3 24.3 15.3
@@
CodeLLaMA (PAL) CodeLLaMA 13B 19.9 39.9 29.9
Platypus-2 (Lee et al., 2023) LLaMA-2 13B 7.1 23.7 15.4
@@
CoSC-Code (Ours) CodeLLaMA 13B 50.3 77.3 63.8 (+1.8)
LLaMA-1 RFT LLaMA-1 34B - 57.9 -
CodeLLaMA (PAL) CodeLLaMA 34B 23.9 53.3 38.6
@@
Baselines. For proprietary models, we present results from an array of SoTA LLMs, such as OpenAI’s ChatGPT (gpt-3.5-turbo), OpenAI’s GPT-4, OpenAI’s GPT-4V, OpenAI’s GPT-4o, Google’s PaLM-2, Google’s Gemini-1.0 Pro, Google’s Gemini-1.0 Ultra, Google’s Gemini-1.5 Flash, Google’s Gemini-1.5 Pro, Anthropic’s Claude-2, Anthropic’s Claude-3 Opus, and Anthropic’s Claude-3.5 Sonnet. By default, we report CoT prompting results, and include PAL (Program-Aided Language model) prompting (Gao et al., 2023) results for selected models. For open-source models, base models comprise LLaMA-2 (Touvron et al., 2023) and CodeLLaMA (Roziere et al., 2023) with CoT and PAL prompting (Gao et al., 2023). Supervised Fine-Tuning (SFT) employs CoT rationales from the original MATH and GSM8K dataset (15k samples) for fine-tuning. Rejection sampling Fine-Tuning (RFT) (Yuan et al., 2023) leverages multiple models to generate diverse reasoning paths for fine-tuning. WizardMath (Luo et al., 2023) augments data using ChatGPT, and conducts SFT and RLHF. Platypus-2 (Lee et al., 2023) is fine-tuned with Open-Platypus reasoning datasets. ToRA (Gou et al., 2023b) uses GPT-4 to generate tool-used trajectories and finetunes on it.
@@ 4.2 Main Results
In our experiments, we observe several findings that demonstrate the effectiveness of our proposed CoSC model on mathematical datasets. We show the experimental results in Table 2. First and foremost, our proposed CoSC consistently outperforms previous state-of-the-art open-source LLMs across all scales. Specifically, our CoSC-Code can achieve an average improvement of 2.6%, 1.8%, and 2.1%, on 7B, 13B, and 34B size, respectively.
@@ 4.3.1 Effect of Each Training Phase
Our CoSC mechanism comprises two training phases: (1) CoSC foundational learning and (2) CoSC self-enhancement. In order to assess the individual contributions of each training phase, we evaluate multiple combinations of the proposed training phases, the results of which are shown in Table 5.
@@
To further enhance the self-correction performance, we utilize the seed model obtained in the first phase to generate additional data without relying on GPT-4. This approach leads to further improvements in performance on the MATH dataset from 42.3% to 47.6% and 47.0% to 50.3%, for 7B and 13B respectively.
@@ 4.3.2 Effect of Multi-Round Reasoning in the Proposed CoSC mechanism
In addition, we conduct a comparison between single-round reasoning and multi-round reasoning using our CoSC mechanism on the test samples from the MATH dataset. The single-round results are obtained by extracting the answer solely from the output of the first round, without any self-correction enabled. To illustrate the impact of multi-round reasoning, we present the comparative results in Table 5. The results clearly demonstrate the effectiveness of multi-round reasoning in rectifying errors and improving overall accuracy. With the 7B model, we observe an improvement in accuracy from 40.2% to 47.6%, while with the 13B model, accuracy increases from 42.4% to 50.3%. These findings highlight the significant benefits of employing multi-round reasoning within our CoSC mechanism. More ablation studies are shown in Appendix D.
@@
Models 7B 13B
#Round=1 #Round=2 #Round=3 #Round=1 #Round=2 #Round=3
CodeLLaMA 100% 0 0 100% 0 0
ToRA-Code 99.9% 0.1% 0 100% 0 0
CoSC-Code (Ours) 78.3% 12.7% 9.0% 79.3% 13.1% 7.6%
@@
          CoSC-Code with multi-round reasoning (Ours)           47.6           50.3
@@ 5 Conclusion
In conclusion, our Chain of Self-Correction (CoSC) mechanism equips Large Language Models (LLMs) with the ability to autonomously validate and refine their outputs. This mechanism facilitates a sequence of self-correction stages that progressively refine the reasoning process, leading to enhanced accuracy in mathematical reasoning. Our two-phase finetuning approach, involving initial training with proprietary model-generated seeding data and subsequent training with self-generated data, effectively enables the implementation of CoSC at a low cost. Through extensive experiments, we have demonstrated the remarkable performance improvement that CoSC brings to traditional mathematical datasets when compared to existing open-source LLMs, even the advanced close-sourced LLMs, such as ChatGPT, GPT-4 and multi-modal LLMs like GPT-4V, Gemini-1.0 Pro, and Gemini-1.0 Ultra. We believe that our CoSC mechanism can provide valuable insights for future research and contribute to the ongoing advancement of LLMs across various domains.
@@ A.2 Prompt for our CoSC on GSM8K
Step 1. Verify the consistency between the question and the code. The "python" code was rewrited, num_days is set to 4, which consistent with the from monday to thursday in "Question".
@@ B.1 Case on MATH testing set
The case for CoSC on MATH testing set is as follows:
@@
Step 1. Verify the consistency between the question and the code. The "python" code was rewrited, equation1 = x - 1 - 7‘ and equation2 = 1 - x - 7‘ were added to solve the two possible equations separately, and then get the sum of all solutions.
@@ B.2 Case on GSM8K testing set
The case for CoSC on GSM8K testing set is as follows:
@@
Appendix C Datasets Details

### v2

@@ 3.2.2 CoSC Self Enhancement
Table 1: Accuracy results (%) on MATH and GSM8K datasets. Vanilla models are tested with CoT. ZS indicates the zero-shot inference without demonstrations. PAL refers to the Program-Aided Language model prompting (Gao et al., 2023). The best results in each section are in bold and the second-best results are underlined.
@@
GPT-4o - -      76.6 (0-shot CoT) 96.1 (8-shot CoT) 86.4
GPT-4V - -      52.9 (4-shot) 92.0 (5-shot CoT) 72.5
GPT-4 (PAL) - -       51.8 (PAL) 94.2 (PAL) 73.0
GPT-4 - -       42.5 (CoT) 92.0 (5-shot CoT) 67.3
ChatGPT (PAL) - -       38.7 (PAL) 78.6 (PAL) 58.7
ChatGPT - -       35.5 (CoT) 80.8 (5-shot CoT) 58.2
Claude-3.5 Sonnet - -       71.1 (0-shot CoT) 96.4 (0-shot CoT) 83.8
Claude-3 Opus - -       60.1 (0-shot CoT) 95.0 (0-shot CoT) 77.6
Gemini-1.5 Pro - -       67.7 (4-shot Minerva) 90.8 (11-shot) 79.3
Gemini-1.5 Flash - -       54.9 (4-shot Minerva) 86.2 (11-shot) 70.6
Gemini-1.0 Ultra - -       53.2 (4-shot Minerva) 88.9 (11-shot) 71.1
@@
CodeLLaMA (PAL) CodeLLaMA 7B      16.6 (PAL) 34.0 (PAL) 25.3
Platypus-2 (Lee et al., 2023) LLaMA-2 7B      5.4 (CoT) 14.4 (CoT) 9.9
@@
CoSC-Code (Ours) CodeLLaMA 7B      47.9 75.1 61.5 (+2.9)
LLaMA-2 LLaMA-2 13B      6.3 (CoT) 24.3 (CoT) 15.3
@@
CodeLLaMA (PAL) CodeLLaMA 13B      19.9 (PAL) 39.9 (PAL) 29.9
Platypus-2 (Lee et al., 2023) LLaMA-2 13B      7.1 (CoT) 23.7 (CoT) 15.4
@@
CoSC-Code (Ours) CodeLLaMA 13B      50.6 77.9 64.3 (+2.3)
CodeLLaMA (PAL) CodeLLaMA 34B      23.9 (PAL) 53.3 (PAL) 38.6
@@
Dense solution sampling. For mathematical questions, there are usually multiple solutions, but using GPT-4 for annotating them is expensive. Therefore, we use the seed model after CoSC foundational learning to resample the questions in the datasets multiple times, which further improves the model generalization ability in answering questions. Specifically, we use the CodeLLaMA-34B model which has the best performance after CoSC foundational learning to perform dense nucleus-sampling on 16k training data questions. Each question in the MATH and GSM8K training datasets is sampled 64 times. We filter out the correct answers based on whether they match the ground-truth.
Dense question sampling. Similarly, there are multiple ways to ask a mathematical question. To improve the model generalization ability of question understanding, we use the data from MetaMath (Yu et al., 2023) to generalize the questions. Each question in MetaMATH is rewritten by simply rephrasing it, self-verifying the conditions in the question, adding if-then questions to reverse the question conditions, etc. Specifically, we use the CodeLLaMA-34B model with CoSC foundational learning to perform nucleus-sampling on the questions, sampling each question once, and filtering out data points based on whether they match the ground-truth.
In the end, we obtain a total of 339k data points, including 37k seeding data generated from GPT-4 and 302k generated from the CodeLLaMA-34B model with CoSC foundational learning. Then, we adopt them to train models from scratch to obtain our final CoSC model.
3.3 Implementation Details
By using 339k data points, including 37k seeding data and 302k self-generated data, we fine-tune the base models of LLaMA-2 (Touvron et al., 2023) and CodeLLaMA (Roziere et al., 2023) to obtain our CoSC and CoSC-Code, respectively. They have different parameter sizes, such as 7B, 13B, and 34B. All models use full-scale fine-tuning. We use the AdamW optimizer with a learning rate of 2e-5 for all models, with a batch size set to 128, training for 1 epoch. To enable training, we use DeepSpeed ZeRO stage 3 (Rajbhandari et al., 2021) and Flash-Attention 2 (Dao, 2023) to optimize the model’s memory usage. During inference, we set a maximum of 3 calls to the Python interpreter and a maximum token length of 2048. The GPT-4 version for CoSC data generation is gpt-4-0613. Our experiments train the models in 7B size with 8 NVIDIA A800 80GB GPUs and train the models in 13B and 34B with 16 NVIDIA A800 80GB GPUs.
4 Experiments
4.1 Evaluation Setup
Datasets. We evaluated models on the most widely used mathematical problem tasks, MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021). The MATH dataset encompasses a total of 12,500 problems, partitioned into 7,500 for training and 5,000 for testing. The GSM8K dataset contains 8,500 problems, with 7,500 for training and 1,000 for testing. These datasets collectively encompass a broad range of mathematical questions, from basic arithmetic to competition level. More details of datasets are provided in Appendix C.1.
Baselines. For proprietary models, we present results from an array of SoTA LLMs, such as OpenAI’s ChatGPT (gpt-3.5-turbo), OpenAI’s GPT-4, OpenAI’s GPT-4V, OpenAI’s GPT-4o, Google’s Gemini-1.0 Ultra, Google’s Gemini-1.5 Flash, Google’s Gemini-1.5 Pro, Anthropic’s Claude-3 Opus, and Anthropic’s Claude-3.5 Sonnet. By default, we report CoT prompting results, and include PAL (Program-Aided Language model) prompting (Gao et al., 2023) results for selected models. For open-source models, base models comprise LLaMA-2 (Touvron et al., 2023) and CodeLLaMA (Roziere et al., 2023) with CoT and PAL prompting (Gao et al., 2023). Supervised Fine-Tuning (SFT) employs CoT rationales from the original MATH and GSM8K dataset (15k samples) for fine-tuning. Rejection sampling Fine-Tuning (RFT) (Yuan et al., 2023) leverages multiple models to generate diverse reasoning paths for fine-tuning. WizardMath (Luo et al., 2023) augments data using ChatGPT, and conducts SFT and RLHF. Platypus-2 (Lee et al., 2023) is fine-tuned with Open-Platypus reasoning datasets. ToRA (Gou et al., 2023b) uses GPT-4 to generate tool-used trajectories and finetunes on it. The prompting methods for each evaluation are shown in Appendix C.2.
@@ 4.2 Main Results
In our experiments, we observe several findings that demonstrate the effectiveness of our proposed CoSC model on mathematical datasets. We show the experimental results in Table 1. First and foremost, our proposed CoSC consistently outperforms previous state-of-the-art open-source LLMs across all scales. Specifically, our CoSC-Code can achieve an average improvement of 2.9%, 2.3%, and 2.1%, on 7B, 13B, and 34B size, respectively.
@@ 4.3.1 Effect of Each Training Phase
Our CoSC mechanism comprises two training phases: (1) CoSC foundational learning and (2) CoSC self-enhancement. In order to assess the individual contributions of each training phase, we evaluate multiple combinations of the proposed training phases, as shown in Table 4.
@@
To further enhance the self-correction performance, we utilize the seed model obtained in the first phase to generate additional data without relying on GPT-4. This approach leads to further improvements in performance on the MATH dataset from 42.3% to 47.9% and 47.0% to 50.6%, for 7B and 13B respectively.
@@ 4.3.2 Effect of Multi-Round Reasoning in the Proposed CoSC mechanism
In addition, we conduct a comparison between single-round reasoning and multi-round reasoning using our CoSC mechanism on the test samples from the MATH dataset. The single-round results are obtained by extracting the answer solely from the output of the first round, without any self-correction enabled. To illustrate the impact of multi-round reasoning, we present the comparative results in Table 4. The results clearly demonstrate the effectiveness of multi-round reasoning in rectifying errors and improving overall accuracy. With the 7B model, we observe an improvement in accuracy from 40.2% to 47.9%, while with the 13B model, accuracy increases from 42.4% to 50.6%. These findings highlight the significant benefits of employing multi-round reasoning within our CoSC mechanism. More ablation studies are shown in Appendix D.
@@
Models Size #Round=1 #Round=2 #Round=3
CodeLLaMA 7B 100% 0 0
13B 100% 0 0
ToRA-Code 7B 100% 0 0
13B 100% 0 0
CoSC-Code (Ours) 7B 78.3% 12.7% 9.0%
13B 79.3% 13.1% 7.6%
@@
CoSC-Code with multi-round reasoning (Ours) 47.9 50.6
@@ 5 Conclusion
In conclusion, our Chain of Self-Correction (CoSC) mechanism equips LLMs with the ability to autonomously validate and refine their outputs. This mechanism facilitates a sequence of self-correction stages that progressively refine the reasoning process, leading to enhanced accuracy in mathematical reasoning. Through extensive experiments, we have demonstrated the remarkable performance improvement that CoSC brings to mathematical datasets. We believe that our CoSC mechanism can provide valuable insights for future research on LLMs across various domains.
Impact Statement
We experiment on two mathematical datasets, including GSM8K and MATH, both of which use MIT License code. The prompts used in these experiments are listed in Appendix A, and we want to emphasize that none of the prompts contain any words that discriminate against any individual or group. Furthermore, prompts would not negatively impact anyone’s safety in this work.
@@ A.2 Prompt for our CoSC on GSM8K
Step 1. Verify the consistency between the question and the code. The "python" code was rewritten, num_days is set to 4, which consistent with the from monday to thursday in "Question".
@@ Appendix B Case of our CoSC on MATH and GSM8K Testing Sets
We would like to show examples how our CoSC model generate a corresponding answer when a question is queried.
@@ B.1 Case on MATH testing set
The case for CoSC on MATH testing set is as follows, where the question is input into our CoSC model and the solution represents the corresponding answer generated from our CoSC model.
@@
Step 1. Verify the consistency between the question and the code. The "python" code was rewritten, equation1 = x - 1 - 7‘ and equation2 = 1 - x - 7‘ were added to solve the two possible equations separately, and then get the sum of all solutions.
@@ B.2 Case on GSM8K testing set
The case for CoSC on GSM8K testing set is as follows, where the question is input into our CoSC model and the solution represents the corresponding answer generated from our CoSC model.
@@
Appendix C Evaluation Setup Details
C.1 Datasets Details
@@
C.2 Prompting Methods for Each Evaluation
Table 1 includes an identifier “ZS”, which denotes whether the LLMs are evaluated in a zero-shot inference setting without demonstrations. To clarify further, we summarize below the prompting methods employed for each evaluation.
Proprietary Models:
  • GPT-4o (OpenAI, 2024): Zero-shot CoT prompting for MATH; 8-shot CoT prompting for GSM8K.
  • GPT-4V (OpenAI, 2023b): 4-shot prompting for MATH; 5-shot CoT prompting for GSM8K.
  • GPT-4 and ChatGPT (OpenAI, 2023a): CoT prompting for MATH; 5-shot CoT prompting for GSM8K.
  • Gemini family (Team et al., 2024): 4-shot Minerva prompting for MATH; 11-shot prompting for GSM8K.
  • Claude family (Anthropic, 2023): Zero-shot CoT prompting for both datasets.
  • PaLM-2 (Anil et al., 2023): 4-shot CoT prompting for MATH; 8-shot CoT prompting for GSM8K.
Open-Source Models:
  • LLaMA-2 (Touvron et al., 2023) and Platypus-2 (Lee et al., 2023): CoT prompting for both datasets.
  • CodeLLaMA (Roziere et al., 2023): Program-Aided Language (PAL) (Gao et al., 2023) prompting for both datasets.
  • LLaMA-2 SFT (Yuan et al., 2023), LLaMA-2 RFT (Yuan et al., 2023), WizardMath (Luo et al., 2023), MetaMath (Yu et al., 2023), ToRA (Gou et al., 2023b), and our CoSC method: Fully zero-shot, requiring no demonstrations.
@@ D.2 Context Lengths for Different Questions
Table 7: Accuracy results (%) on MATH and GSM8K datasets of ToRA-Code-7B, ToRA-Code-7B under CoSC prompts and our CoSC-Code-7B.
       Models        MATH        GSM8K        AVG
       ToRA-Code-7B        44.6        72.6        58.6
       ToRA-Code-7B under CoSC prompts        42.8        68.0        55.4
       CoSC-Code-7B (Ours)        47.9        75.1        61.5
@@ D.3 Alleviated Issues by our CoSC
D.4 Results of ToRA under CoSC prompts
To explore whether only using CoSC prompts can embed self-correction ability in LLMs, we conduct an experiment using the CoSC prompts for evaluation on ToRA-Code-7B. The CoSC prompts are same as those used in CoSC seeding data generation in Appendix A. The results on the MATH and GSM8K datasets are shown in Table 7.
As shown in Table 7, applying CoSC prompting to ToRA not only fails to outperform the original ToRA model but also results in a decline in performance. As demonstrated in Table 4, ToRA inherently lacks the robust multi-round reasoning capabilities needed for effective self-correction. When CoSC prompting is applied, it introduces complexity that the model is ill-equipped to handle, leading to confusion and errors in the iterative process. Similarly, during the development of the CoSC algorithm, we also attempt to apply self-correction prompts to the base CodeLLaMA model. However, this approach did not yield good performance and was significantly lower than the previous state-of-the-art results in open-source models. This led us to adopt a fine-tuning strategy instead.
In contrast, our CoSC model, which integrates self-correction as an inherent capability via fine-tuning, achieves superior results on both datasets. These findings suggest that for open-source LLMs, few-shot prompting alone is insufficient to effectively enable self-correction. The lack of significant gains from prompting further underscores the limitations of relying solely on in-context examples. Therefore, we argue that embedding self-correction as an inherent capability through fine-tuning is essential for truly endowing LLMs with robust self-correction abilities.
Moreover, by integrating self-correction directly into the training process, our approach allows models to perform self-correction autonomously in a zero-shot setting during inference, eliminating the need for external feedback or few-shot demonstrations. This self-correction mechanism enables even weaker LLMs to achieve significant improvements in mathematical reasoning—enhancements that are unattainable through prompting methods alone. Additionally, our CoSC framework is open-source, making these advancements accessible to the broader research community. We believe this represents a pivotal step toward democratizing advanced reasoning capabilities and fostering further innovation.